# Large Language Models powered Network Attack Detection: Architecture, Opportunities and Case Study

### v1

@@ III-B LLM as Encoder
In network attack detection, LLM can also act as a representation learner, the Encoder. It transforms network raw traffic into a vector representation, i.e., tokens or embeddings. This vector encapsulates the essential characteristics of the network data, making it easier to process and classify further. LLMs can not only learn explicit features but also mine latent features that humans cannot understand, such as in-context correlations of network traffic data.
As shown in Fig. 3, the typical pipeline of LLM as Encoder based network attack detection consists of three parts: Tokenizer, LLM Encoder, and MLP classifier.
@@
LLM Encoder: The tokens are fed into the LLM Encoder. The LLM Encoder, based on the pre-trained LLM architecture, processes the tokens. It uses the learned patterns and knowledge from its pre-training to generate a vector representation for the input tokens. Usually two types of LLM model can be used in network threat detection, BERT of bidirectional encoder, and GPT of auto-regressive embeddings generation.
MLP classifier: By adding a output layer to a pretrained LLM, significant classification results can be obtained in network threat detection. Usually a simple MLP (Multi-Layer Perceptron) is used as the output layer. The MLP classifier takes the vector representation output by the LLM Encoder and uses it to classify the network traffic as benign or malicious.
@@
Generalization: An LLM Encoder trained or fine-tuned on a specific set of network traffic may not be able to effectively generate accurate vector representations for networks with different traffic pattern or attack vectors. It’s a critical challenge for LLM encoder to extract features against data drifts.
Scalability: As the scale of network traffic and the number of network devices continue to grow exponentially, the scalability of LLM-based encoders becomes a pressing issue. Scaling up the LLM encoder to handle this massive data volume requires not only more computational resources but also efficient distributed computing strategies.
@@ III-C LLM as Predictor
As the classical anomaly detection methods, LLMs could act as Predictor to detect network anomalies. The LLMs, being trained with normal or baseline traffic during its pre-training or fine-tuning phase, predict the most likely next or masked token in a network traffic sequence. If the actual traffic deviates from the predicted tokens, it could indicate a threat.
Fig. 4 demonstrate the pipeline of LLM-as-Predictor. The language model of BERT and GPT can be used to execute anomaly detection. For BERT model, the masked token is predicted by bi-direction contexture sequence. For GPT model, the next token in the sequence is predicted in a auto-regressive way. LLM trained or fine-tuned only with benign traffic data are expected to exhibit significant errors and low predictive probability when they encounter the context of abnormal traffic during the test.
As in FIg. 4, the LLM-based anomaly detection framework consists of three parts:
Tokenizer: As new network traffic is captured, tokenize it and then mask parts of tokens for predictions. For BERT model, the key tokens in the sequence are masked. For GPT model, the tokens are fed into LLM one-by-one.
Fine-tuned LLM: Use the normal network traffic data (tokenized and labeled) to pre-training or fine-tune the LLM. The fine-tuning process adjusts the pre-trained weights of the LLM to better fit the characteristics of the network traffic data. This involves training the LLM to predict the next token or masked token in a normal traffic sequence.
Anomaly Detection: The prediction error is calculated by cross-entropy loss that occurs between the predicted token and the input traffic. Additionally, the predictive probability is a value with the highest probability of token that can appear in the prediction. When the probability of a predicted token is low or the prediction error is large, the respective traffic traffic is considered rare to find in the normal context and is identified as an anomaly.
LLM-as-predictor explore the LLMs’ capability on token generations. It has a crucial assumption that Abnormal traffic’s in-context pattern is different from that of normal traffic. It faces several challenges in network attack detection.
@@
Pattern Drift: Same as classical anomaly-based detection, the distribution of network traffic data changes continuously over time. This requires the model to adapt to these changes in a timely manner. Otherwise, the probability scores calculated by the model based on historical data may not accurately reflect the current abnormal situations, resulting in a large number of false positives or false negatives.
@@ IV Case Study: A Design of LLM-based DDos Detection
In this section, we introduce a system design as a case study that detects Carpet Bombing DDoS Attacks (CBA) by LLMs and achieves high accuracy, especially zero-shot capability.
IV-A LLM-based Encoder and Flow-Correlation Mining
This system aims to detect a new type of DDoS, Carpet Bombing DDos Attacks (CBA), on the backbone network of telecom operators [14]. CBA is currently an increased threat that targets large-scale network infrastructure such as public cloud, internet data centers, and even ISPs. It simultaneously floods a group of IPs (e.g., a few of C-subset networks) , and congests their shared bottleneck link. Since the end-host is not the explicit target but the bottleneck link is, the legacy end-host based detection method is not applicable anymore [15].
Leveraging the potential of LLM as Encoder, the proposed LLM-based system mines the multi-flows contextual correlations to enhance the detection features. Since the attackers during a CBA usually use the same attack tools, there exist strong correlations among malicious flows. Hence, we can use LLMs’ self-attention mechanism to capture the correlations between tokens (words) in sequence.
The framework consist of four parts: Flow Sequentializer, Flow Tokenizer, LLM Encoder and Classifier. It organizes network flows into a flow sequence by Flow Sequentializer, and projects each flow into the semantic space of LLMs by Flow Tokenizer. LLM acts as encoder to transforming the token into flow embedding. The inter-flow correlations are explored by the LLMs backbone. Finally, the detection task is modeled as a binary classification task, and the flow will be classified based on the flow embeddings outputs by the LLMs backbone.
In this detection pipeline, it deals with two challenges flow sequence organization and the modality alignment between flow data and text.
Flow Sequentializer. It designs novel methods for flow sequence organization. Since the position encoding of LLMs, the position of each flow in a sequence is important. However, due to the huge number of users in the backbone network and the network jitter, the flows will be randomly arrived. Organizing flow sequence by the arrived order straight-forwardly will affect the correlation extracting of LLMs. Therefore, The purpose Flow Sequentializer to generate flow sequence by steps of Flow Sorting, Equal Frequency Binning, and Vertical Flows Assembling. According to the steps, the flow in the same position will have the same characteristics, making the pattern of the flow sequence easy to learn.
Flow Tokenizer. To transfer LLMs’ ability to capture inter-token correlations to inter-flow correlation mining, it treats flow statistics as a new modality and aligns them to the semantic space of LLMs through alignment. To achieve this, it designs a learnable Flow Tokenizer. Flow Tokenizer encoders each flow in the flow sequences to a high-dimension token embedding. And inter-flow correlations will be explored by a frozen LLM backbone. The output embeddings will be classified by the Classification Projection, and output the binary label of malicious or benign.
@@ IV-B Evaluation
The system is implemented on a Linux server with an AMD Ryzen-9 7950X 16-core CPU and an NVIDIA GeForce RTX-4090 GPU, and uses Llama2-7b-chat as the backbone, configuring the model’s floating-point precision to bfloat16 to enable GPU execution. It is evaluated in both simulation and actual world trace. In the simulation test, a multi-vector dataset and a zero-shot dataset are made by the benchmark CIC-DDoS2019 and MAWI dataset. In trace based test, a NetFlow dataset is collected from a Top-3 national-wide operator.
It shows the best detection performance across all scenarios. Especially, in zero-shot scenarios, it’s F1 score improved by upto $24\%$ compared to other methods, indicating it’s potential to detect new types of attacks. In real-world NetFlow trace, it improved by at least $35.1\%$, showcasing its significant advantage in real-world scenarios. The test results for the above two scenarios are shown in Fig. 5.
@@ V Conclusion
With the explosively break-down of LLM, Applications of LLM in network security has developed rapidly in recent years. This paper studies LLM-powered network attack detection, from a novel perspective of what roles LLMs can function in network attack detection. We present a holistic view of the architecture of LLM-powered network attack detection, including unsupervised pre-training, supervised fine-tuning, inference and detection. Research opportunities and challenges are next discussed. Finally, we present our case study on LLM-based carpet bombing DDoS detection, showcasing the innovative design. The propose framework successfully achieve good accuracy for malicious traffic detection. Through evaluation, we have demonstrated the remarkable effectiveness of our approach, especially for real-world traffic, surpassing existing system with a nearly 35% improvement.

### v2

@@ III-B LLM as Encoder
In malicious traffic detection, LLM can also act as a representation learner, the Encoder. It transforms network raw traffic into a vector representation, i.e., tokens or embeddings. This vector encapsulates the essential characteristics of the network data, making it easier to process and classify further. LLMs can not only learn explicit features but also mine latent features that humans cannot understand, such as in-context correlations of network traffic data.
As shown in Fig. 3, the typical pipeline of LLM as Encoder based network threat detection consists of three parts: Tokenizer, LLM Encoder, and MLP classifier.
@@
LLM Encoder: The tokens are fed into the LLM Encoder. The LLM Encoder, based on the pre-trained LLM architecture, processes the tokens. It uses the learned patterns and knowledge from its pre-training to generate a vector representation for the input tokens. Usually two types of LLM model can be used in malicious traffic detection, BERT of bidirectional encoder, and GPT of auto-regressive embeddings generation.
Given a sequence of input tokens $w_{i}$, the encoded representations can be presented as:
$h_{1},h_{2},...,h_{T}=Encoder(w_{1},...,w_{T})$ (1)
Here $h_{i}$ is the output embeddings which is used for classification in the following steps.
MLP classifier: By adding a output layer to a pretrained LLM, traffic classification results can be obtained. Usually a simple MLP (Multi-Layer Perceptron) is used as the output layer. The MLP classifier takes the vector representation output by the LLM Encoder and uses it to classify the network traffic as benign or malicious.
@@
Generalization: An LLM Encoder trained or fine-tuned on a specific set of network traffic may not be able to effectively generate accurate vector representations for networks with different traffic pattern or attack vectors. It’s a critical challenge for LLM encoder to extract features against traffic drifts.
Scalability: As the scale of network traffic and the number of network devices continue to grow exponentially, the scalability of LLM-based encoders becomes a critical issue. Scaling up the LLM encoder to handle this massive data volume requires not only more computational resources but also efficient distributed computing.
@@ III-C LLM as Predictor
As the classical anomaly detection methods, LLMs could act as Predictor to detect network traffic anomalies. The LLMs, being trained with normal or baseline traffic during its pre-training or fine-tuning phase, predict the most likely next or masked token in a network traffic sequence. If the actual traffic deviates from the predicted tokens, it could indicate a anomaly.
Fig. 4 demonstrate the pipeline of LLM-as-Predictor. The language model of BERT and GPT can be used to execute anomaly detection. For BERT model, the masked token is predicted by bi-direction contexture sequence. For GPT model, the next token in the sequence is predicted in a auto-regressive way. LLM is trained or fine-tuned only with benign traffic data. It is expected to exhibit significant errors and low probability when they encounter abnormal traffic during the test.
As in Fig. 4, the framework of LLM-based anomaly detection consists of three parts:
Tokenizer: As network traffic is captured, it tokenizes them and then masks parts of tokens for predictions. For BERT model, the key tokens in the sequence are masked. For GPT model, the tokens are fed into LLM one-by-one.
Fine-tuned LLM: Use the normal network traffic data (tokenized and labeled) to fine-tune the LLM. The fine-tuning process adjusts the pre-trained weights of the LLM to better fit the characteristics of the normal network traffic. This involves training the LLM to predict the next token or masked token in a normal traffic sequence.
Anomaly Detection: In the detection, the prediction error is calculated by cross-entropy loss that occurs between the predicted token and the input traffic. Additionally, the predictive probability is a value with the highest probability of token that can appear in the prediction. When the probability of a predicted token is low or the prediction error is large, the respective traffic traffic is considered rare to find in the normal context and is identified as an anomaly.
LLM-as-predictor explore the LLMs’ capability on token generations. It has a crucial assumption that abnormal traffic’s in-context pattern is different from that of normal traffic. It faces several challenges:
@@
Pattern Drift: As in classical anomaly-based detection, the distribution of network traffic data changes continuously over time. This requires the model to adapt to these changes in a timely manner. Otherwise, the probability scores calculated by the model based on historical data may not accurately reflect the current abnormal situations, resulting in a large number of false positives or false negatives.
@@ IV Case Study: A Design of LLM-based DDos Detection
In this section, we introduce a system design as a case study that LLMs detect Carpet Bombing DDoS Attacks (CBA) with high accuracy and especially zero-shot capability.
IV-A LLM-based Encoder in Flow-Correlation Mining
This system is design to detect a typical network attack, Carpet Bombing DDos Attacks (CBA), on the backbone network of telecom operators [14]. CBA is currently an increased threat that targets large-scale network infrastructure such as public cloud, internet data centers, and ISP networks. It simultaneously floods a group of IPs, and overwhelms their shared bottleneck link. Since the end-host is not the explicit target but the bottleneck link is, the legacy victim-based detection method is not applicable anymore [15].
Leveraging the potential of LLM as Encoder, the proposed LLM-based system mines the multi-flows contextual correlations to enhance the detection features. Since the attackers during a CBA usually use the same attack tools, there exist strong correlations among malicious flows. Hence, we can use LLMs’ self-attention mechanism to capture the correlations between tokens (network flows) in sequence.
The framework consist of four parts: Flow Sequentializer, Flow Tokenizer, LLM Encoder and Classifier. It organizes network flows into a flow sequence by Flow Sequentializer, and projects each flow into the semantic space of LLMs by Flow Tokenizer. LLM acts as encoder to transforming the token into flow embedding. The inter-flow correlations are explored by the LLMs backbone. Finally, the detection task is modeled as a binary classification task, and the flow will be classified based on the flow embeddings outputs by the LLMs.
In the pipeline, it deals with two challenges, flow-sequence construction and modality alignment between flows and text.
Flow Sequentializer. Since the position encoding of LLMs, the position of each flow in a sequence is important. However, due to the huge number of users in the backbone network and the network jitter, the flows will be randomly arrived. Therefore, The purpose Flow Sequentializer is to generate flow sequence by steps of Flow Sorting, Equal Frequency Binning, and Vertical Flows Assembling. As a result, the flow in the same position will have the same characteristics, making the pattern of the flow sequence easy to learn.
Flow Tokenizer. For inter-flow correlation mining, it treats network flow as a new modality and aligns them to the semantic space of LLMs. It designs a learnable Flow Tokenizer which encodes flow sequences into a high-dimension token embedding. And inter-flow correlations will be explored by LLM. The embedding representations will be classified by a MLP, and output the binary label of malicious or benign.
@@ IV-B Evaluation
The system is implemented on a Linux server with an AMD Ryzen-9 7950X 16-core CPU and an NVIDIA GeForce RTX-4090 GPU, and uses Llama2-7b-chat as the backbone, configuring the model’s floating-point precision to bfloat16.
DoLLM is evaluated in both simulation and real-world trace. In the simulation test, we mix the CIC-DDoS2019 and MAWI datasets to create two datasets, each with a 1:1 ratio of malicious to benign flows. The first dataset includes three attack vectors: DNS, NTP, and SYN. The second dataset consists of the remaining eight vectors from CIC-DDoS2019. We train on the former and test on the latter to evaluate zero-shot capability. In the real-world test, a NetFlow dataset is collected from a Top-3 nationwide operator, and the ratio of malicious to benign flows is 1:10. We compare the performance of DoLLM with four ML-based methods, including the supervised learning method XGBoost, the unsupervised anomaly detection method AutoEncoder, and two sequence-based models: LSTM and Transformer. These methods encompass the majority of current data-driven detection mechanisms. We use the F1-score as the evaluation metric.
It shows the best detection performance across all scenarios. Especially, in zero-shot scenarios, its F1 score improved by up to $24\%$ compared to other methods, indicating its potential to detect new types of attacks. In real-world NetFlow trace, it improved by at least $35.1\%$ compared to other methods, showcasing its significant advantage in real-world scenarios. The test results for the above two scenarios are shown in Fig. 5.
@@ V Conclusion
With the explosively break-down of LLM, Applications of LLM in network security has developed rapidly in recent years. This paper studies LLM-powered malicious traffic detection, from a novel perspective of what roles LLMs can function. We present a holistic view of the architecture of LLM-powered malicious traffic detection, including unsupervised pre-training, supervised fine-tuning and ICL, inference and detection. Research opportunities and challenges are next discussed. Finally, we present our case study on LLM-based carpet bombing DDoS detection, showcasing the innovative design. The proposed framework successfully achieve good accuracy for malicious traffic detection, surpassing existing system with a nearly 35% improvement.
The proposed LLMs architecture is a general framework for classification tasks. It can be not only applied to malicious traffic detection, but also to any other classification task. LLMs can act as the Classifier, Encoder, and Predictor to mine massive data, uncover the latent features and representations, and achieve the accuracy classification. With the remarkable capabilities of context-understanding and commonsense knowledge, LLMs would provide a new paradigm for traffic classification and network threat detections. In the future, we would continue exploring the potentials of LLMs in more network tasks, such as abnormal log analysis, malware detection, encrypted traffic classification, and beyond.